OpenStudy (anonymous):

9) A gas at 400K and 1atm has it's pressure doubled. What is the new temperature if the amount of gas and volume of gas remains the same?

OpenStudy (abmon98):

For a fixed mass of gas at Volume, the Pressure is directly proportional to the kelvin temperature. P1/P2=T1/T2 1/2=400/T2

OpenStudy (anonymous):

Thanks @Abmon98